# Paley-Wiener Theorems For Slice Regular Functions

### v1

@@ 1. Introduction
In this paper we will prove the compact and non-compact types of Paley-Wiener theorem for slice regular functions. To state the main results, we introduce some notations. Denote by $\mathbb{H}$ the quaternions space, by $\mathscr{R}_{l}(U)$ the class of left slice regular functions on the open set $U\subset\mathbb{H},$ by $\mathbb{H}_{+}$ the open right half-space of $\mathbb{H},$ by $H^{p}(\mathbb{H}_{+})$ the Hardy space on $\mathbb{H}_{+},$ and by $\mathbb{S}$ the unit sphere of purely imaginary quaternions.
@@
$f\left(x+Iy\right)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{0}e^{\left(x+Iy\right)t}\mathscr{F}_{E}\left(f\right)\left(t\right)\mathrm{d}t,$
where $\mathscr{F}_{E}$ is defined in Definition 2.16.
Main Result II: Sufficient condition for non-compact type. Let $F:\mathbb{R}\to\mathbb{H}$ and $F\in L^{p}(\mathbb{R})$ for $1\leq p\leq 2$. Assume that there exists a quaternionic imaginary unit $I\in\mathbb{S}$ such that $\operatorname{supp}\mathscr{F}_{I}\left(F\right)\subset\left(-\infty,0\right]$. Then for every $J\in\mathbb{S}$, and $x>0,y\in\mathbb{R},$
is in $H^{p}\left(\mathbb{H}_{+}\right)$, where $\mathscr{F}_{E}\left(f\right)=\mathscr{F}_{I}\left(F\right)$.
@@ 2. Preliminaries
In fact, we note that Definition 2.1 is a re-elaboration of a definition given by Gentili, Stoppato and Struppa (see [19, Chapter 1]).
@@
Proposition 2.3 ([6] Representation Formula).
@@
are two real-valued functions (see [22, Proposition 10] and [7, Proposition 2.5]). if this is the case and if $x+Iy\in U$, it follows at once that $\Re f\left(x+Iy\right)=\alpha\left(x,y\right)$ and that $\lvert f\left(x+Iy\right)\rvert=\sqrt{\alpha\left(x,y\right)^{2}+\beta\left(x,y\right)^{2}}$, whence $\lvert f\left(x+Iy\right)\rvert=\lvert f\left(x+Jy\right)\rvert$ for all $I,J\in\mathbb{S}$. Moreover, the following result is given.
@@
Let $f\in\mathscr{N}_{l}(\mathbb{H}_{+})\cap H^{1}(\mathbb{H}_{+})$. Then for every $I,J\in\mathbb{S}$ and $t\in\mathbb{R}$, we have
@@
    $\overline{\mathscr{F}_{I}\left(F^{I}\right)(t)}=\mathscr{F}_{I}\left(F^{I}\right)(t)$;
@@
    $\mathscr{F}_{J}\left(F^{J}\right)(t)=\mathscr{F}_{I}\left(F^{I}\right)(t)$.
@@
Since $f\in\mathscr{N}_{l}(\mathbb{H}_{+})$, we have $f(\overline{\mathbf{q}})=\overline{f(\mathbf{q})}$ for $\mathbf{q}\in\mathbb{H}_{+}$. For every $I\in\mathbb{S}$, let $F^{I}(y)=f(Iy)$ be the NTBL of $f$ at $Iy$. Then we have $F^{I}(-y)=\overline{F^{I}(y)}$. By $f\in H^{1}(\mathbb{H}_{+})$, we obtain that $F^{I}\in L^{1}(\mathbb{R})$ and
@@
It follows that
$\overline{\mathscr{F}_{I}(F^{I})(t)}=$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{Iyt}\overline{F^{I}(y)}\mathrm{d}y$
$=$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{Iyt}F^{I}(-y)\mathrm{d}y$
$=$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{-Iyt}F^{I}(y)\mathrm{d}y$
$=$ $\mathscr{F}_{I}(F^{I})(t).$
Hence, we can get that $\mathscr{F}_{I}(F^{I})$ is real-valued. Note that there exist two real-valued functions $\alpha(0,y)$ and $\beta(0,y)$ such that
$F^{I}(y)=\alpha(0,y)+I\beta(0,y).$
Then
$\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\left[\cos(yt)\beta(0,y)-\sin(yt)\alpha(0,y)\right]\mathrm{d}y=0$
and
$\mathscr{F}_{I}(F^{I})(t)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\left[\cos(yt)\alpha(0,y)+\sin(yt)\beta(0,y)\right]\mathrm{d}y.$
For every $J\in\mathbb{S}$, let $F^{J}(y)$ be that NTBL of $f$ at $Jy$. It is easy to show that
$F^{J}(y)=\alpha(0,y)+J\beta(0,y).$
and $|F^{J}(y)|^{2}=|F^{I}(y)|^{2}$. Thus, we can obtain that $F^{J}\in L^{1}(\mathbb{R})$ and
$\mathscr{F}_{J}(F^{J})(t)=$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{-Jyt}F^{J}(y)\mathrm{d}y$
$=$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\left[\cos(yt)\beta(0,y)-\sin(yt)\alpha(0,y)\right]\mathrm{d}y$
$+\frac{J}{\sqrt{2\pi}}\int_{\mathbb{R}}\left[\cos(yt)\alpha(0,y)+\sin(yt)\beta(0,y)\right]\mathrm{d}y$
$=$ $\mathscr{F}_{I}(F^{I})(t).$
@@
where $\left\{1,I,J,IJ\right\}$ is a basis of $\mathbb{H}_{+}$. It is easy to see that if $f\in\mathscr{N}_{l}(\mathbb{H}_{+})\cap H^{1}(\mathbb{H}_{+})$ then $f_{m}\in\mathscr{N}_{l}(\mathbb{H}_{+})\cap H^{1}(\mathbb{H}_{+})$, $0\leq m\leq 3$. It follows that
@@ 3.1. Compact type Paley-Wiener theorem
Based on the argument in [33, Theorem 19.3] (see also [28]), we can immediately obtain Lemma 3.2. For completeness, we include a brief proof.
@@
Suppose $f\in\mathscr{N}_{l}\left(\mathbb{H}\right)$ with $f|_{\mathbb{R}}\in L^{2}(\mathbb{R})$, and $A,C>0$ are two positive numbers. Assume that there exists a quaternionic imaginary unit $I\in\mathbb{S}$, such that $|f(x+Iy)|\leq Ce^{A|x+Iy|}$ for all $x,y\in\mathbb{R}$. Then for every $J\in\mathbb{S}$ and $x,y\in\mathbb{R}$, $\operatorname{supp}\mathscr{F}_{J}\left(f|_{\mathbb{R}}\right)\subset\left[-A,A\right]$ and
@@
For a fixed $I\in\mathbb{S}$, by $f\in\mathscr{N}_{l}\left(\mathbb{H}\right)$, we have $f\left(\mathbb{C}_{I}\right)\subseteq\mathbb{C}_{I}$ and $\overline{f\left(x-Iy\right)}=f\left(x+Iy\right)$. Since $|f(x+Iy)|\leq Ce^{A|x+Iy|}$ for all $x,y\in\mathbb{R}$ and $f|_{\mathbb{R}}\in L^{2}(\mathbb{R})$, by Lemma 3.2 we have $\operatorname{supp}\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\subset\left[-A,A\right]$ and
(3.6) $f\left(x+Iy\right)=\frac{1}{\sqrt{2\pi}}\int_{-A}^{A}e^{I\left(x+Iy\right)t}\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\left(t\right)\mathrm{d}t.$
By Proposition 2.11 and Plancherel’s theorem, we can conclude that $\mathscr{F}_{I}(f|_{\mathbb{R}})$ is $\mathbb{C}_{I}$-valued satisfying
(3.7) $\overline{\mathscr{F}_{I}(f|_{\mathbb{R}})(t)}=\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)$
and
(3.8) $\mathscr{F}_{J}(f|_{\mathbb{R}})(t)=\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]+\frac{JI}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(t)\right]$
for all $t\in\mathbb{R}$ and $J\in\mathbb{S}$. Thus, we have $f\left(x+Iy\right)=\alpha\left(x,y\right)+I\beta\left(x,y\right)$ with
$\alpha\left(x,y\right)$
$=$ $\frac{1}{2}\left[f(x+Iy)+f(x-Iy)\right]$
$=$ $\frac{1}{2\sqrt{2\pi}}\int_{-A}^{A}e^{Ixt}\left(e^{yt}+e^{-yt}\right)\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\left(t\right)\mathrm{d}t$
$=$ $\frac{1}{2\sqrt{2\pi}}\int_{-A}^{A}\left(e^{yt}+e^{-yt}\right)\Big{\{}\cos\left(xt\right)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\ \ -\sin\left(xt\right)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t$
$+\frac{I}{2\sqrt{2\pi}}\int_{-A}^{A}\left(e^{yt}+e^{-yt}\right)\Big{\{}\cos(xt)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\quad\quad+\sin(xt)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t$
and
$\beta\left(x,y\right)$
$=$ $\frac{-I}{2}\left[f(x+Iy)-f(x-Iy)\right]$
$=$ $\frac{-I}{2\sqrt{2\pi}}\int_{-A}^{A}e^{Ixt}\left(e^{yt}-e^{-yt}\right)\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\left(t\right)\mathrm{d}t$
$=$ $\frac{-I}{2\sqrt{2\pi}}\int_{-A}^{A}\left(e^{yt}-e^{-yt}\right)\Big{\{}\cos\left(xt\right)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\ \ -\sin\left(xt\right)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t$
$+\frac{1}{2\sqrt{2\pi}}\int_{-A}^{A}\left(e^{yt}-e^{-yt}\right)\Big{\{}\cos(xt)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\quad\quad+\sin(xt)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t.$
Using the fact that $\alpha(x,y)$ and $\beta(x,y)$ are real-valued functions, we have
$\frac{I}{2\sqrt{2\pi}}\int_{-A}^{A}\left(e^{yt}+e^{-yt}\right)\Big{\{}\cos(xt)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\quad\quad+\sin(xt)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t=0$
and
$\frac{-I}{2\sqrt{2\pi}}\int_{-A}^{A}\left(e^{yt}-e^{-yt}\right)\Big{\{}\cos\left(xt\right)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\ \ -\sin\left(xt\right)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t=0.$
Therefore,
$\alpha(x,y)=$ $\frac{1}{\sqrt{2\pi}}\int_{-A}^{A}e^{-yt}\Big{\{}\cos\left(xt\right)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\ \ -\sin\left(xt\right)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t$
and
$\beta(x,y)=$ $\frac{1}{\sqrt{2\pi}}\int_{-A}^{A}e^{-yt}\Big{\{}\cos(xt)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\quad\quad+\sin(xt)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t.$
It follows that
$f\left(x+Jy\right)$
$=$ $\alpha(x,y)+J\beta(x,y)$
$=$ $\frac{1}{\sqrt{2\pi}}\int_{-A}^{A}e^{-yt}\Big{\{}\cos\left(xt\right)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\ \ -\sin\left(xt\right)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t$
$+J\frac{1}{\sqrt{2\pi}}\int_{-A}^{A}e^{-yt}\Big{\{}\cos(xt)\frac{-I}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad\quad\quad+\sin(xt)\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]\Big{\}}\mathrm{d}t$
$=$ $\frac{1}{\sqrt{2\pi}}\int_{-A}^{A}e^{-yt}e^{Jxt}\Big{\{}\frac{1}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(t)+\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)\right]$
$\quad\quad\quad\quad+\frac{JI}{2}\left[\mathscr{F}_{I}(f|_{\mathbb{R}})(-t)-\mathscr{F}_{I}(f|_{\mathbb{R}})(t)\right]\Big{\}}\mathrm{d}t$
$=$ $\frac{1}{\sqrt{2\pi}}\int_{-A}^{A}e^{J\left(x+Jy\right)t}\mathscr{F}_{J}\left(f|_{\mathbb{R}}\right)\left(t\right)\mathrm{d}t.$
Moreover, (3.7) and (3.8) imply that $\lvert\mathscr{F}_{J}\left(f|_{\mathbb{R}}\right)(t)\rvert^{2}=\lvert\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)(t)\rvert^{2}$ holds for all $t\in\mathbb{R}$. Thus, we have $\operatorname{supp}\mathscr{F}_{J}\left(f|_{\mathbb{R}}\right)=\operatorname{supp}\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\subset\left[-A,A\right]$. ∎
Lemma 3.4.
Let $f\in\mathscr{N}_{l}\left(\mathbb{H}\right)$ with $f|_{\mathbb{R}}\in L^{2}(\mathbb{R})$, and $A>0$ be a positive number. Assume that there exists a quaternionic imaginary unit $I\in\mathbb{S}$, such that $\operatorname{supp}\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\subset\left[-A,A\right]$. Then for every $J\in\mathbb{S}$ and $x,y\in\mathbb{R}$, $|f(x+Jy)|\leq Ce^{A|x+Jy|}$ with $C>0$.
Proof.
For a fixed $I\in\mathbb{S}$, let
$\Psi(x+Iy)=\frac{1}{\sqrt{2\pi}}\int_{-A}^{A}e^{I\left(x+Iy\right)t}\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\left(t\right)\mathrm{d}t.$
[…]
$\lvert\Psi(x+Iy)\rvert$ $\leq\frac{1}{\sqrt{2\pi}}\int_{-A}^{A}\lvert e^{I\left(x+Iy\right)t}\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\left(t\right)\rvert\mathrm{d}t$
$\leq\frac{1}{\sqrt{2\pi}}\left(\int_{-A}^{A}\lvert e^{I\left(x+Iy\right)t}\rvert^{2}\mathrm{d}t\right)^{\frac{1}{2}}\left(\int_{-A}^{A}\lvert\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\left(t\right)\rvert^{2}\mathrm{d}t\right)^{\frac{1}{2}}$
$\leq\frac{\sqrt{2A}}{\sqrt{2\pi}}e^{A|y|}\lVert f|_{\mathbb{R}}\rVert_{L^{2}(\mathbb{R})}=Ce^{A|x+Iy|}.$
Since $f\in\mathscr{N}_{l}\left(\mathbb{H}\right)$, we have $f\left(\mathbb{C}_{I}\right)\subset\mathbb{C}_{I}$ and $\overline{\partial_{I}}f\left(x+Iy\right)=0.$ Using the fact that $\Psi(x)=f(x)$ and both functions are holomorphic in $\mathbb{C}_{I}$, we conclude that $f\left(x+Iy\right)=\Psi(x+Iy)$. Hence, for every $J\in\mathbb{S}$, we have $\lvert f\left(x+Jy\right)\rvert=\lvert f\left(x+Iy\right)\rvert\leq Ce^{A|x+Iy|}=Ce^{A|x+Jy|}$.

### v2

@@ 1. Introduction
In this paper we will prove the compact and non-compact types of Paley-Wiener theorem for slice regular functions. To state the main results, we introduce some notations. Denote by $\mathbb{H}$ the quaternions space, by $\mathscr{R}_{l}(U)$ the class of left slice regular functions on the open set $U\subset\mathbb{H},$ by $\mathbb{H}_{+}$ the open right half-space of $\mathbb{H},$ by $H^{p}(\mathbb{H}_{+})$ the Hardy space on $\mathbb{H}_{+},$ by $\mathscr{F}_{I}$ the one-dimensional left-sided quaternion Fourier transform (see Definition 2.12), by $\mathscr{F}_{E}$ the essential Fourier transform (see Definition 2.18), and by $\mathbb{S}$ the unit sphere of purely imaginary quaternions.
@@
Main Result II: Sufficient condition for non-compact type. Let $F:\mathbb{R}\to\mathbb{H}$ and $F\in L^{p}(\mathbb{R})$ for $1\leq p\leq 2$. Assume that there exists a quaternionic imaginary unit $I\in\mathbb{S}$ such that $\operatorname{supp}\mathscr{F}_{I}\left(F\right)\subset\left(-\infty,0\right]$. Then there exists a function $f\in H^{p}(\mathbb{H}_{+})$, whose non tangential boundary boundary limit at $Iy$ is $F(y)$, whence $\mathscr{F}_{E}\left(f\right)=\mathscr{F}_{I}\left(F\right)$ and
$f\left(x+Jy\right):=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{0}e^{\left(x+Jy\right)t}\mathscr{F}_{I}\left(F\right)\left(t\right)\mathrm{d}t,$
for all $x>0,y\in\mathbb{R}$, $J\in\mathbb{S}$.
@@ 2. Preliminaries
In fact, we note that Definition 2.1 is a re-elaboration of a definition given by Gentili, Stoppato and Struppa (see [20, Chapter 1]).
@@
Proposition 2.3 ([21]).
Let $B(0,R)$ be a ball centered at the origin and of radius $R$. A function $f\in\mathscr{R}_{l}(B(0,R))$ if, and only if, it has a series expansion of the form
$f(\mathbf{q})=\sum_{n=0}^{\infty}\mathbf{q}^{n}\frac{1}{n!}\frac{\partial^{n}f}{\partial x^{n}}(0)$
converging on $B(0,R)$. In particular if $f\in\mathscr{R}_{l}(B(0,R))$ then it is $C^{\infty}$ on $B(0,R)$.
Proposition 2.4 ([20, 21]).
Let $f\in\mathscr{R}_{l}(B(0,R))$ on the ball $B(0,R)$ centered at the origin with radius $R$, and let
$f(\mathbf{q})=\sum_{n=0}^{\infty}\mathbf{q}^{n}a_{n}$
be power series expansion. Then $f\in\mathscr{N}_{l}(B(0,R))$ if, and only if, $\{a_{n}\}_{n\in\mathbb{N}}\subset\mathbb{R}$.
Proposition 2.5 ([7, 24] Representation Formula).
@@
are two real-valued functions (see [23, Proposition 10] and [8, Proposition 2.5]). if this is the case and if $x+Iy\in U$, it follows at once that $\Re f\left(x+Iy\right)=\alpha\left(x,y\right)$ and that $\lvert f\left(x+Iy\right)\rvert=\sqrt{\alpha\left(x,y\right)^{2}+\beta\left(x,y\right)^{2}}$, whence $\lvert f\left(x+Iy\right)\rvert=\lvert f\left(x+Jy\right)\rvert$ for all $I,J\in\mathbb{S}$. Moreover, the following result is given.
@@
Let $f\in\mathscr{N}_{l}(\mathbb{H}_{+})\cap H^{1}(\mathbb{H}_{+})$. Then for all $I,J\in\mathbb{S}$, the function $\mathscr{F}_{I}(F^{I})$ is real-valued and coincides with $\mathscr{F}_{J}(F^{J})$, i.e.,
@@
    $\overline{\mathscr{F}_{I}(F^{I})(t)}=\mathscr{F}_{I}(F^{I})(t)$;
@@
    $\mathscr{F}_{J}(F^{J})(t)=\mathscr{F}_{I}(F^{I})(t).$
@@
Since $f\in\mathscr{N}_{l}(\mathbb{H}_{+})\cap H^{1}(\mathbb{H}_{+})$, for every $I\in\mathbb{S}$, we obtain that $F^{I}\in L^{1}(\mathbb{R})$ and
$F^{I}(y)=\alpha(0,y)+I\beta(0,y),$
where $\alpha(0,y)$ and $\beta(0,y)$ are two real-valued functions. It follows that
@@
Note that
$\mathscr{F}_{I}(F^{I})(t)=$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\left[\cos(yt)\alpha(0,y)+\sin(yt)\beta(0,y)\right]\mathrm{d}y$
$+\frac{I}{\sqrt{2\pi}}\int_{\mathbb{R}}\left[\cos(yt)\beta(0,y)-\sin(yt)\alpha(0,y)\right]\mathrm{d}y.$
The second integral vanishes because, for $t$ fixed, the integrand is odd. Hence, for all $J\in\mathbb{S}$, we have $\mathscr{F}_{J}(F^{J})=\mathscr{F}_{I}(F^{I})$ which is real-valued. ∎
@@
where $I,J$ are mutually orthogonal quaternionic imaginary units. It is easy to see that if $f\in H^{1}(\mathbb{H}_{+})$ then $f_{m}\in\mathscr{N}_{l}(\mathbb{H}_{+})\cap H^{1}(\mathbb{H}_{+})$, $0\leq m\leq 3$. It follows that
@@ 3.1. Compact type Paley-Wiener theorem
Based on the argument in [36, Theorem 19.3] (see also [30]), we can immediately obtain Lemma 3.2. For completeness, we include a brief proof.
@@
Suppose $f\in\mathscr{N}_{l}\left(\mathbb{H}\right)$ with $f|_{\mathbb{R}}\in L^{2}(\mathbb{R})$, and let $A,C>0$ be two positive constants. Then we have the following results:
  1. (i)
    Assume that there exists a quaternionic imaginary unit $I\in\mathbb{S}$, such that $\operatorname{supp}\mathscr{F}_{I}\left(f|_{\mathbb{R}}\right)\subset\left[-A,A\right]$. Then for every $J\in\mathbb{S}$ and $x,y\in\mathbb{R}$, $|f(x+Jy)|\leq Ce^{A|x+Jy|}$.
  2. (ii)
    Assume that there exists a quaternionic imaginary unit $I\in\mathbb{S}$ such that $|f(x+Iy)|\leq Ce^{A|x+Iy|}$ for all $x,y\in\mathbb{R}$. Then, for every $J\in\mathbb{S}$ and $x,y\in\mathbb{R}$, $\operatorname{supp}\mathscr{F}_{J}\left(f|_{\mathbb{R}}\right)\subset\left[-A,A\right]$ and
@@
For a fixed $I\in\mathbb{S}$, since $f\in\mathscr{N}_{l}(\mathbb{H})$, we have $|f(x+Jy)|=|f(x+Iy)|$ for any $J\in\mathbb{S}$. Using Lemma 3.2, we immediately get (i). Moverover, if $\mathscr{F}_{I}(f|_{\mathbb{R}})$ is supported in $[-A,A]$ then so is $\mathscr{F}_{J}(f|_{\mathbb{R}})$ for any $J\in\mathbb{S}$, thanks to property (3) in Proposition 2.13. Thus, Lemma 3.2 implies (ii). ∎